zepdrix (zepdrix):

Not really a question. Just thinking about a calculus concept. Comment if you like. :)

zepdrix (zepdrix):

Here are some common derivative rules that we know. For these I'm assuming f=f(x), and a is constant. \[\Large\rm y=f^a,\qquad y'=a(f^{a-1})f'\qquad\qquad \text{Power Rule}\]\[\Large\rm y=a^f,\qquad y'=a^f (f') \ln a\qquad\qquad \text{Exponential Rule}\] I'm just trying to think about what happens when we have a function to a function and take it's derivative and then see if I can come up with some generalization. You guys can add insight if you want, I just want to show my thought process. So for some functions f=f(x), g=g(x),\[\Large\rm y=f^g\]log of each side,\[\Large\rm \ln y=g \ln f\]Taking derivative,\[\Large\rm \frac{1}{y}y'=g' \ln f+g \frac{1}{f}f'\]Multiply y to the other side,\[\Large\rm y'=y\left[g' \ln f+g \frac{1}{f}f'\right]\]plug in the y, get a common denominator in the brackets,\[\Large\rm y'=f^g\left[\frac{f g' \ln f+f' g}{f}\right]\]Let's ummm, pull the 1/f out and combine it with the term in front,\[\Large\rm y'=f^{g-1}\left[f g' \ln f+f' g\right]\]Then let's distribute the thing into the brackets,\[\Large\rm y'=f^g(g')\ln f+g(f^{g-1})f'\]Now that I'm looking at it, I can see that getting the common denominator and pulling out the 1/f was maybe a couple of unnecessary steps hehe. Anyway, check this out, this is kind of neat. \[\Large\rm y'=\text{Exponential Rule + Power Rule}\]We started out using product rule and a bunch of fun stuff, and it looks like we end up with the sum of these two rules that we listed at the start. Hmm interestinggggg.

zepdrix (zepdrix):

Maybe this is already detailed on wikipedia or somewhere, I dunno. I just wanted to work through it and try to make some kind of connection :p

OpenStudy (fibonaccichick666):

I think it works minus the y=exponential rule +power rule part

zepdrix (zepdrix):

No? :o It looks like it matches those rules.. hmm

OpenStudy (fibonaccichick666):

it's a combination of em, not a straight addition of the two

OpenStudy (fibonaccichick666):

nope i retract. I was reading product rule for some reason

zepdrix (zepdrix):

So let's try using that on a simple example problem maybe, \[\Large\rm y=x^{2x}\]without using log of each side and all that business.\[\Large\rm y=\text{Exponential Rule + Power Rule}\]\[\Large\rm y'=x^{2x}(2x)'\ln x+2x(x^{2x-1})x'\]\[\Large\rm y'=2x^{2x}\ln x+2x(x^{2x-1})\]\[\Large\rm y'=2x^{2x}\left[\ln x+1\right]\]Which is the correct answer. Hmm that's kinda neat :)

OpenStudy (fibonaccichick666):

yea, nice derivation

zepdrix (zepdrix):

Maybe I'm being a bit sloppy calling it "exponential rule" when our base is clearly x, but we're treating it as constant for that step, i dunno

OpenStudy (fibonaccichick666):

I assume there are some limits to the function though

OpenStudy (anonymous):

..... So much math *dies* I can't take it.... all the big words... ahhhhhh

zepdrix (zepdrix):

XD

OpenStudy (fibonaccichick666):

might be a only works for positive f type of thing... I'm not sure

zepdrix (zepdrix):

ya probably, something like this ends up being crazy anyway,\[\Large\rm (-2)^x\]even without taking its derivative hehe

OpenStudy (fibonaccichick666):

shall we find out?

zepdrix (zepdrix):

:x

OpenStudy (fibonaccichick666):

Oh wolfram, how I love thee http://www.wolframalpha.com/input/?i=%28%28-2%29%5Ex%29%27

OpenStudy (fibonaccichick666):

so, have to dip into complex analysis to pull that bad boy off

zepdrix (zepdrix):

Ohh interesting

zepdrix (zepdrix):

\[\Large\rm y=(-2)^x\]\[\Large\rm y'=(-2)^x \ln(-2)\qquad\qquad \text{Exponential Rule}\]\[\Large\rm y'=(-2)^x\left[\ln(2)+\ln(-1)\right]\]And yah gotta dip into the complex here.. hmm trying to remember >.<

OpenStudy (fibonaccichick666):

i pi

zepdrix (zepdrix):

Oh oh oh right!\[\Large\rm e^{i \pi}=-1\]\[\Large\rm \ln e^{i \pi}=\ln(-1)\]\[\Large\rm i \pi=\ln(-1)\]

zepdrix (zepdrix):

Hmm neato :)

OpenStudy (fibonaccichick666):

yea, cool trick

OpenStudy (fibonaccichick666):

wanna check my theory? I may need to prove it and get published if it works

zepdrix (zepdrix):

maybeeee <.< I'm feeling kinda lazy right now lol

OpenStudy (fibonaccichick666):

haha, well, I tagged ya. It's all about coloring :D

OpenStudy (kainui):

Interesting, I have always noticed that there was this thing \[\large \frac{d}{dx} x^x = xx^{x-1}+x^x \ln x\] but I never really thought much past that, i was to distracted while taking calculus 2 haha.

OpenStudy (kainui):

Ok let's see if we can take this a step further, what about this? \[\Large y=f^{g^h}\]

zepdrix (zepdrix):

buhaha >.> of course YOU would do something like that lol

OpenStudy (fibonaccichick666):

... well, question cause I'm not sure, isn't that the same as \(lny=hglnf\)?

OpenStudy (fibonaccichick666):

so it's just a triple iteration of the product rule?

OpenStudy (kainui):

Nah it's more like \[\ln( \ln y)=h \ln g + \ln(\ln f)\]

zepdrix (zepdrix):

\[\Large\rm \ln y=g^h \ln f\]If we're allowed to use our first result, then I don't think we need to log again.

OpenStudy (kainui):

Oh true

zepdrix (zepdrix):

But if you want to start from scratch, yah it's a bit of work :)

OpenStudy (fibonaccichick666):

so that's where I'm wondering cause we'd have \[lny=ln f^{g^h}\]

OpenStudy (fibonaccichick666):

so we can bring the exponents out front right?

OpenStudy (fibonaccichick666):

but can we pull both...I actually don't know

zepdrix (zepdrix):

Fib, I think you were confusing \(\Large\rm \left(f^g\right)^h\) with \(\Large\rm f^{(g^h)}\). We want the second one :D

OpenStudy (fibonaccichick666):

oh

OpenStudy (fibonaccichick666):

yep

OpenStudy (kainui):

So I guess this is getting gross but applied twice: \[\large y' = (g^h)f^{g^h-1}f'+(g^h)'f^{g^h} \\ \large y' = (g^h)f^{g^h-1}f'+(hg^{h-1}g'+g^hh')f^{g^h}\]

OpenStudy (fibonaccichick666):

I agree with kainui now

OpenStudy (kainui):

So I guess this is getting gross but applied twice: \[\large y =f ^{g^h} \\ \large y' = (g^h)f^{g^h-1}f'+(g^h)'f^{g^h} \\ \large y' = (g^h)f^{g^h-1}f'+(hg^{h-1}g'+g^hh')f^{g^h}\]

OpenStudy (fibonaccichick666):

it seems recursive

zepdrix (zepdrix):

Ooo interesting :3

zepdrix (zepdrix):

Uh oh the site is tweaking out on me :( Too much LaTeX in this thread maybe lol

OpenStudy (fibonaccichick666):

lol potentially

OpenStudy (kainui):

Yeah so we could keep doing this for \[\large y=f^{g^{h^\cdots}}\] So now what if we did it infinitely, what would the series be of terms in the derivative?

OpenStudy (fibonaccichick666):

...That would be a serious series...

OpenStudy (dan815):

did u figure out the pattern for nth exponent functions of functions

OpenStudy (kainui):

Maybe we should index them \[\Large y=f_1^{f_2^{f_3^\cdots}}\]

OpenStudy (dan815):

ya lol that looks a lot neater

OpenStudy (dan815):

the recursion is always coming from the exponential rule side, i think its better to just see the pattern with lns but i guess that wud defeat the purpose

OpenStudy (dan815):

we need new ntoation for mulitple exponents

OpenStudy (dan815):